# Invariant Subspace Decomposition

### v1

@@ 2.1 Invariant and residual subspaces
We now construct the two linear subspaces $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ that allow us to express the true time-varying parameter as the solution to two separate optimizations as in (5). For a linear subspace $\mathcal{S}\subseteq\mathbb{R}^{p}$, denote by $\Pi_{\mathcal{S}}$ the orthogonal projection matrix from $\mathbb{R}^{p}$ onto $\mathcal{S}$. We call a collection of pairwise orthogonal linear subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{q}\subseteq\mathbb{R}^{p}$ with $\bigoplus_{j=1}^{q}\mathcal{S}_{j}=\mathbb{R}^{p}$ an orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ (of cardinality $q$) if for all $i,j\in\{1,\dots,q\}$ with $i\neq j$, and for all $t\in[n]$ it holds that
@@
Moreover, we call an orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ irreducible if there is no other orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ with strictly larger cardinality.
For an orthogonal (not necessarily irreducible) partition $\{\mathcal{S}_{j}\}_{j=1}^{q}$ w.r.t. $(X_{t})_{t\in[n]}$ we show in Lemma 3 that
$\gamma_{0,t}=\operatorname{Var}(X_{t})^{-1}\operatorname{Cov}(X_{t},Y_{t})=\sum_{j=1}^{q}\operatorname{Var}\left(\Pi_{\mathcal{S}_{j}}X_{t}\right)^{\dagger}\operatorname{Cov}\left(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t}\right),$ (7)
where $(\cdot)^{\dagger}$ denotes the Moore-Penrose pseudoinverse. A direct computation (see Lemma 4) further shows that, for all $j\in\{1,\ldots,q\}$ and all $t\in[n]$,
$\operatorname*{arg\,max}_{\beta\in\mathcal{S}_{j}}\operatorname{\Delta Var}_{t}(\beta)=\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})^{\dagger}\operatorname{Cov}(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t}),$ (8)
where to show that the argmax is unique we used that $\operatorname{Var}(X_{t})$ is invertible by assumption. Combining (7) and (8) implies that
$\gamma_{0,t}=\sum_{j=1}^{q}\operatorname*{arg\,max}_{\beta\in\mathcal{S}_{j}}\operatorname{\Delta Var}_{t}(\beta).$
Therefore, any orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ allows to split the optimization (3) of the explained variance into separate optimizations over the individual subspaces. In order to leverage all available observations in at least some of the optimizations by pooling the explained variance as in (5), we need the optimizer to remain constant over time. More formally, we call a linear subspace $\mathcal{S}\subseteq\mathbb{R}^{p}$ opt-invariant on $[n]$ if for all $t,s\in[n]$ it holds that
$\operatorname*{arg\,max}_{\beta\in\mathcal{S}}\operatorname{\Delta Var}_{t}(\beta)=\operatorname*{arg\,max}_{\beta\in\mathcal{S}}\operatorname{\Delta Var}_{s}(\beta),$ (9)
namely the optimizer of $\operatorname{\Delta Var}_{t}(\beta)$ over $\beta\in\mathcal{S}$ remains invariant across $t\in[n]$. For an irreducible orthogonal partition $\{\mathcal{S}_{j}\}_{j=1}^{q}$, we now define the invariant subspace $\mathcal{S}^{\operatorname{inv}}$ and the residual subspace $\mathcal{S}^{\operatorname{res}}$ by
@@
respectively. The invariant subspace $\mathcal{S}^{\operatorname{inv}}$ is opt-invariant (see Lemma 5). Moreover, by definition, it holds that $\mathcal{S}^{\operatorname{res}}=\left(\mathcal{S}^{\operatorname{inv}}\right)^{\perp}$, where $(\cdot)^{\perp}$ denotes the orthogonal complement in $\mathbb{R}^{p}$, and that $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ is an orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ (see Lemma 6). To ensure that these two spaces do not depend on the chosen irreducible subspace partition we introduce the following assumption.
@@
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ and $\{\bar{\mathcal{S}}_{j}\}_{j=1}^{\bar{q}}$ be two irreducible orthogonal partitions w.r.t. $(X_{t})_{t\in[n]}$. Then, it holds that
@@
Assumption 1 is for example satisfied if an irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ is unique but, as we show in Appendix  A.1, such uniqueness does not necessarily hold and, furthermore, Assumption 1 may be violated. Assumption 1 is, however, a mild assumption; for example, the uniqueness of an irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ is satisfied if there exists at least one $t\in[n]$ such that all eigenvalues of $\Sigma_{t}$, which we have assumed to be non-zero, are distinct (see Lemma 7 and Proposition 1 below). Whenever Assumption 1 holds, the invariant and residual spaces $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ do not depend on which irreducible orthogonal partition is used in their construction.
@@
where $\sigma_{1,t}$ and $\sigma_{2,t}$ are two fixed sequences sampled as two independent i.i.d. samples from a uniform distribution on $[0,1]$. In this example, we have that an irreducible orthogonal partition is given by the two spaces
@@
which does not depend on $t$ (it can be verified that the same does not hold for $\Pi_{\mathcal{S}_{1}}\gamma_{0,t}$). As can be seen from the proof of Proposition 2 (i) below, it holds that
@@
For all irreducible orthogonal partitions w.r.t. $(X)_{t\in[n]}$, $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ defined in (10) satisfy the following two conditions: (i) for all $t\in\mathbb{N}$ it holds that $\operatorname{Cov}(\Pi_{\mathcal{S}^{\operatorname{inv}}}X_{t},\Pi_{\mathcal{S}^{\operatorname{res}}}X_{t})=0$ and (ii) $\mathcal{S}^{\operatorname{inv}}$ is opt-invariant on $\mathbb{N}$.
As shown in the following theorem, Assumption 2 ensures that the two sets satisfy a separation of the form (5) as desired. For unobserved time points $t\in\mathbb{N}\setminus[n]$, Assumption 2 does not require (6) to hold for all $i,j\in\{1,\dots,q\}$, but only for all $i,j\in\{1,\dots,q\}$ such that $\mathcal{S}_{i}\subseteq\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}_{j}\subseteq\mathcal{S}^{\operatorname{res}}$.
@@
Assume Assumption 2 is satisfied. Let $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ be defined as in (10) for an arbitrary irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$. Then, it holds for all $t\in\mathbb{N}$ that
@@
where $\overline{\operatorname{\Delta Var}}(\beta)\coloneqq\frac{1}{n}\sum_{t=1}^{n}\operatorname{\Delta Var}_{t}(\beta)$. Moreover, if Assumption 1 is satisfied, the separation in (11) is independent of the considered irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$.
@@ 2.1.1 Identifying invariant and residual subspaces using joint block diagonalization
Let $U\in\mathbb{R}^{p\times p}$ be an irreducible joint block diagonalizer of $(\Sigma_{t})_{t\in[n]}$. For all $j\in\{1,\dots,q^{U}_{\max}\}$, let $S_{j}\subseteq\{1,\dots,p\}$ denote the subset of indices corresponding to the $j$-th common block $\tilde{\Sigma}_{t,j}$. Moreover, let $u^{k}$ denote the $k$-th column of $U$ and, for all $j\in\{1,\dots,q^{U}_{\max}\}$, define
@@
Then, $\{\mathcal{S}_{j}\}_{j=1}^{q^{U}_{\max}}$ is an irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$.
The reverse of Proposition 1 is true as well, as shown in Lemma 2. From Proposition 1 it also follows for all $j\in\{1,\dots,q^{U}_{\max}\}$ that the orthogonal projection matrix onto the subspace $\mathcal{S}_{j}$ can be expressed as $\Pi_{\mathcal{S}_{j}}=U^{S_{j}}(U^{S_{j}})^{\top}$, where we denote by $U^{S_{j}}$ the submatrix of $U$ formed by the columns indexed by $S_{j}$. Moreover, if Assumption 1 is satisfied, any irreducible joint block diagonalizer $U$, via its corresponding irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ constructed in Proposition 1, leads to the same (unique) invariant and residual subspaces defined in (10).
It is clear that Assumption 1 is automatically satisfied whenever the joint block diagonalization is unique up to trivial indeterminacies, that is, if for all orthogonal matrices $U,U^{\prime}\in\mathbb{R}^{p\times p}$ that jointly block diagonalize the set $(\Sigma_{t})_{t\in[n]}$ into $q^{U}_{\max}$ irreducible common blocks, it holds that $U$ is equal to $U^{\prime}$ up to block permutations and block-wise isometric transformations. Explicit conditions under which uniqueness of joint block diagonalization is satisfied can be found, for example, in the works by De Lathauwer (2008); Murota et al. (2010).
@@ 2.2 Invariant component
The parameter $\beta^{\operatorname{inv}}$ corresponds to the pooled OLS solution obtained by regressing $Y_{t}$ on the projected predictors $\Pi_{\mathcal{S}^{\operatorname{inv}}}X_{t}$, and can be computed using all observations in $[n]$. The procedure to find $\beta^{\operatorname{inv}}$ by first identifying $\mathcal{S}^{\operatorname{inv}}$ via joint block diagonalization and then using Proposition 2 $(i)$ below is later summarized in Algorithm 1; the last step (line $10$) is based on Definition 2, (7) and (8). Proposition 2 summarizes some of the properties of $\beta^{\operatorname{inv}}$.
@@ 2.4 Population ISD algorithm
The full ISD procedure is summarized in Algorithm 1. In the algorithm, the invariant and residual subspaces are identified as described at the end of Section 2.1.1. If Assumption 1 is not satisfied, then the decomposition in (16) and therefore the output of Algorithm 1 depends on the irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ used.
In Appendix B we show that a decomposition of the true time-varying parameter similar to (16) is also obtained when considering partitions w.r.t. $(X_{t})_{t\in[n]}$ that are not orthogonal, i.e., such that (6) holds but the subspaces in the partition are not necessarily pairwise orthogonal. In this case, in particular, we can still find an invariant and residual parameter as in (13) and (15), respectively, where the matrix $U$ is now a non-orthogonal irreducible joint block diagonalizer.
@@
5:$U,\{S_{j}\}_{j=1}^{q}\leftarrow\operatorname{irreducibleJointBlockDiagonalizer}((\Sigma_{t})_{t\in[n]})$
6:$S^{\operatorname{inv}},S^{\operatorname{res}}\leftarrow\emptyset$
7:for $j=1,\dots,q$ do
8:     $\Pi_{\mathcal{S}_{j}}\leftarrow U^{S_{j}}(U^{S_{j}})^{\top}$
9:     if $\Pi_{\mathcal{S}_{j}}\gamma_{0,t}$ constant in $[n]$  then $S^{\operatorname{inv}}\leftarrow S^{\operatorname{inv}}\cup S_{j}$
@@
11:$\Pi_{\mathcal{S}^{\operatorname{inv}}}\leftarrow U^{\operatorname{inv}}(U^{\operatorname{inv}})^{\top}$
@@
3 ISD tasks: Zero-shot generalization and time adaptation
We now formally introduce the two tasks that can be solved with the ISD framework introduced in the previous section. Assume we observe $n$ observations $(X_{t},Y_{t})_{t\in[n]}$ from model (1), which we call historical data. Furthermore, assume we observe a second set of observations $(X_{t},Y_{t})_{t\in\mathcal{I}^{\operatorname{ad}}}$ from the same model but succeeding the historical data, which we call adaptation data. Formally, we assume the adaptation window $\mathcal{I}^{\operatorname{ad}}\subseteq\mathbb{N}$ is an interval of consecutive time points with $m:=|\mathcal{I}^{\operatorname{ad}}|>p$ and $\min_{t\in\mathcal{I}^{\operatorname{ad}}}t>n$. Finally, denote by $t^{*}>\max_{t\in\mathcal{I}^{\operatorname{ad}}}$ a time point of interest that occurs after the adaptation data and assume that for all $t\in\mathcal{I}^{\operatorname{ad}}\cup t^{*}$, the quantities $\gamma_{0,t}$, $\operatorname{Var}(X_{t})$ and $\operatorname{Var}(\epsilon_{t})$ in model (1) are constant (in practice, being approximately constant is sufficient).
Our goal is now to predict $Y_{t^{*}}$ from $X_{t^{*}}$. We distinguish two possible prediction tasks: (i) the zero-shot task, in which we assume that we only have access to the historical data and (ii) the adaptation task, in which we assume that we additionally have access to the adaptation data. The full setup including the different tasks is visualized in Figure 3. Throughout the remainder of this section we assume that the invariant and residual subspaces are known (they can be computed from the joint distributions), which simplifies the theoretical analysis.

### v2

@@ 2.1 Invariant and residual subspaces
We now construct the two linear subspaces $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ that allow us to express the true time-varying parameter as the solution to two separate optimizations as in (3). For a linear subspace $\mathcal{S}\subseteq\mathbb{R}^{p}$, denote by $\Pi_{\mathcal{S}}$$\in\mathbb{R}^{p\times p}$ the orthogonal projection matrix from $\mathbb{R}^{p}$ onto $\mathcal{S}$, that is, a symmetric matrix such that $\Pi_{\mathcal{S}}^{2}=\Pi_{\mathcal{S}}$ and, for all vectors $v\in\mathbb{R}^{p}$, $\Pi_{\mathcal{S}}v\in\mathcal{S}$ and $(v-\Pi_{\mathcal{S}}v)^{\top}\Pi_{\mathcal{S}}v=0$. We call a collection of pairwise orthogonal linear subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{q}\subseteq\mathbb{R}^{p}$ with $\bigoplus_{j=1}^{q}\mathcal{S}_{j}=\mathbb{R}^{p}$ an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition (of cardinality $q$) if for all $i,j\in\{1,\dots,q\}$ with $i\neq j$, and for all $t\in[n]$ it holds that
@@
Moreover, we define an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition as irreducible if there is no orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition with strictly larger cardinality. We will construct $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ from an irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition.
Given a (not necessarily irreducible) orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition $\{\mathcal{S}_{j}\}_{j=1}^{q}$ the true time-varying parameter can be expressed as the sum of $q$ orthogonal components each lying in one subspace of the partition. These components can be expressed in terms of the covariates and of the response at time $t$ as shown by the following lemma.
Lemma 1.
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition. Then it holds for all $t\in[n]$ that $\gamma_{0,t}=\sum_{j=1}^{q}\Pi_{\mathcal{S}_{j}}\gamma_{0,t}$ and for all $j\in\{1,\dots,q\}$ that
$\Pi_{\mathcal{S}_{j}}\gamma_{0,t}=\operatorname{Var}\left(\Pi_{\mathcal{S}_{j}}X_{t}\right)^{\dagger}\operatorname{Cov}\left(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t}\right)$ (7)
where $(\cdot)^{\dagger}$ denotes the Moore-Penrose pseudoinverse.
Each component is in particular the projection of $\gamma_{0,t}$ on the corresponding subspace $\mathcal{S}_{j}$, as well as a maximizer of the explained variance in $\mathcal{S}_{j}$, as shown by Lemma 2.
Lemma 2.
Let $\mathcal{N}\subseteq\mathbb{N}$ and let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an orthogonal and $(X_{t})_{t\in\mathcal{N}}$-decorrelating partition.222 This is defined analogously to an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition. Then it holds for all $j\in\{1,\dots,q\}$ and for all $t\in\mathcal{N}$ that
$\operatorname*{arg\,max}_{\beta\in\mathcal{S}_{j}}\operatorname{\Delta Var}_{t}(\beta)=\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})^{\dagger}\operatorname{Cov}(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t}).$ (8)
Combining Lemmas 1 and 2 implies that
$\gamma_{0,t}=\sum_{j=1}^{q}\operatorname*{arg\,max}_{\beta\in\mathcal{S}_{j}}\operatorname{\Delta Var}_{t}(\beta).$ (9)
Therefore, any orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition allows us to split the optimization (2) of the explained variance into separate optimizations over the individual subspaces. In order to leverage all available observations in at least some of the optimizations by pooling the explained variance as in (3), we need the optimizer to remain constant over time. More formally, we call a linear subspace $\mathcal{S}\subseteq\mathbb{R}^{p}$ opt-invariant (optimum invariant) on $[n]$ if for all $t,s\in[n]$ it holds that
$\operatorname*{arg\,max}_{\beta\in\mathcal{S}}\operatorname{\Delta Var}_{t}(\beta)=\operatorname*{arg\,max}_{\beta\in\mathcal{S}}\operatorname{\Delta Var}_{s}(\beta).$ (10)
For all terms in (9) corresponding to an opt-invariant subspace, we can—by definition of opt-invariance—use all time-points in the optimization. For an irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition $\{\mathcal{S}_{j}\}_{j=1}^{q}$, we therefore define the invariant subspace $\mathcal{S}^{\operatorname{inv}}$ and the residual subspace $\mathcal{S}^{\operatorname{res}}$ by
@@
respectively. The invariant subspace $\mathcal{S}^{\operatorname{inv}}$ is opt-invariant (see Lemma 7). Moreover, by definition, it holds that $\mathcal{S}^{\operatorname{res}}=\left(\mathcal{S}^{\operatorname{inv}}\right)^{\perp}$, where $(\cdot)^{\perp}$ denotes the orthogonal complement in $\mathbb{R}^{p}$, and that $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ is an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition (see Lemma 6). To ensure that these two spaces do not depend on the chosen irreducible orthogonal partition, we introduce the following assumption.
@@
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ and $\{\bar{\mathcal{S}}_{j}\}_{j=1}^{\bar{q}}$ be two irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partitions. Then, it holds that
@@
Assumption 1 is for example satisfied if an irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition is unique. As we show in Appendix  A.1, the uniqueness does not always hold (e.g., if for all $t\in[n]$ the multiplicity of some eigenvalues of $\Sigma_{t}$ is larger than one and shared across all such matrices). Assumption 1 is, however, a mild assumption; for example, the uniqueness of an irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition is satisfied if there exists at least one $t\in[n]$ such that all eigenvalues of $\Sigma_{t}$, which we have assumed to be non-zero, are distinct (see Lemma 8 and Proposition 1 below). Whenever Assumption 1 holds, the invariant and residual spaces $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ do not depend on which irreducible orthogonal partition is used in their construction.
@@
where $\sigma_{1,t}$ and $\sigma_{2,t}$ are two fixed sequences sampled as two independent i.i.d. samples from a uniform distribution on $[0,1]$. In this example, we have that an irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition is given by the two spaces
@@
which does not depend on $t$ (it can be verified that the same does not hold for $\Pi_{\mathcal{S}_{1}}\gamma_{0,t}$). As shown in Lemmas 1 and 2, it holds that
@@
For all irreducible orthogonal and $(X)_{t\in[n]}$-decorrelating partitions, $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ defined in (11) satisfy the following two conditions: (i) for all $t\in\mathbb{N}$ it holds that $\operatorname{Cov}(\Pi_{\mathcal{S}^{\operatorname{inv}}}X_{t},\Pi_{\mathcal{S}^{\operatorname{res}}}X_{t})=0$ and (ii) $\mathcal{S}^{\operatorname{inv}}$ is opt-invariant on $\mathbb{N}$.
As shown in the following theorem, Assumption 2 ensures that the two sets satisfy a separation of the form (3) for all observed and unobserved time points $t\in\mathbb{N}$, as desired. For unobserved time points $t\in\mathbb{N}\setminus[n]$, Assumption 2 does not require (6) to hold for all $i,j\in\{1,\dots,q\}$, but only for all $i,j\in\{1,\dots,q\}$ such that $\mathcal{S}_{i}\subseteq\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}_{j}\subseteq\mathcal{S}^{\operatorname{res}}$.
@@
Assume Assumption 2 is satisfied. Let $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ be defined as in (11) for an arbitrary irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition. Then, it holds for all $t\in\mathbb{N}$ that
@@
where $\overline{\operatorname{\Delta Var}}(\beta)\coloneqq\frac{1}{n}\sum_{t=1}^{n}\operatorname{\Delta Var}_{t}(\beta)$. Moreover, if Assumption 1 is satisfied, the separation in (12) is independent of the considered irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition.
The proof of Theorem 1 can be found in Appendix E.1 and relies on the fact that, under Assumption 2, the invariant and residual subspace form an orthogonal and $(X_{t})_{t\in\mathbb{N}}$-decorrelating partition.
@@ 2.1.1 Identifying invariant and residual subspaces using joint block diagonalization
  1. (i)
    Let $U\in\mathbb{R}^{p\times p}$ be a joint block diagonalizer of $(\Sigma_{t})_{t\in[n]}$. For all $j\in\{1,\dots,q^{U}_{\max}\}$, let $S_{j}\subseteq\{1,\dots,p\}$ denote the subset of indices corresponding to the $j$-th common block $\tilde{\Sigma}_{t,j}$. Moreover, let $u^{k}$ denote the $k$-th column of $U$ and, for all $j\in\{1,\dots,q^{U}_{\max}\}$, define
@@
    Then, $\{\mathcal{S}_{j}\}_{j=1}^{q^{U}_{\max}}$ is an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition. Moreover, if the joint block diagonalizer is irreducible, then the corresponding orthogonal partition is irreducible.
  2. (ii)
    The converse is also true. Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition. Then, there exists a joint block diagonalizer $U\in\mathbb{R}^{p\times p}$ of $(\Sigma_{t})_{t\in[n]}$ such that for all $t\in[n]$ the matrix $\tilde{\Sigma}_{t}\coloneqq U^{\top}\Sigma_{t}U$ is block diagonal with $q$ diagonal blocks $\widetilde{\Sigma}_{t,j}=(U^{S_{j}})^{\top}\Sigma_{t}U^{S_{j}}$, $j\in\{1,\dots,q\}$ of dimension $|S_{j}|=\operatorname{dim}(\mathcal{S}_{j})$, where $S_{j}\subseteq\{1,\dots,p\}$ indexes a subset of the columns of $U$. If the orthogonal partition is irreducible, then $U$ is an irreducible joint block diagonalizer. Moreover, $\Pi_{\mathcal{S}_{j}}=U^{S_{j}}(U^{S_{j}})^{\top}$.
If Assumption 1 is satisfied, any irreducible joint block diagonalizer $U$, via its corresponding irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition constructed in Proposition 1(i), leads to the same (unique) invariant and residual subspaces defined in (11).
It is clear that Assumption 1 is automatically satisfied whenever the joint block diagonalization is unique up to trivial indeterminacies, that is, if for all orthogonal matrices $U,U^{\prime}\in\mathbb{R}^{p\times p}$ that jointly block diagonalize the set $(\Sigma_{t})_{t\in[n]}$ into $q^{U}_{\max}$ irreducible common blocks, it holds that $U$ is equal to $U^{\prime}$ up to block permutations and block-wise isometric transformations. Explicit conditions under which uniqueness of joint block diagonalization is satisfied can be found, for example, in the works by De Lathauwer (2008); Murota et al. (2010). Intuitively, these conditions are satisfied whenever there is sufficient variability across time in the covariance matrices $(\Sigma_{t})_{t\in[n]}$.
@@ 2.2 Invariant component
The parameter $\beta^{\operatorname{inv}}$ corresponds to the pooled OLS solution obtained by regressing $Y_{t}$ on the projected predictors $\Pi_{\mathcal{S}^{\operatorname{inv}}}X_{t}$, and can be computed using all observations in $[n]$. The whole procedure to find $\beta^{\operatorname{inv}}$, by first identifying $\mathcal{S}^{\operatorname{inv}}$ via joint block diagonalization and then using Proposition 2 $(i)$ below, is summarized in Algorithm 1 (see Section 2.4). Proposition 2 summarizes some of the properties of $\beta^{\operatorname{inv}}$.
@@ 2.4 Population ISD algorithm
The full ISD procedure is summarized in Algorithm 1. In the algorithm, the invariant and residual subspaces are identified through joint block diagoanlization as described at the end of Section 2.1.1. The number of subspaces $q$ in the irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition is inferred by the joint block diagonalization algorithm. If Assumption 1 is not satisfied, then the decomposition in (17) and therefore the output of Algorithm 1 depends on the irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition used.
In Appendix C we show that a decomposition of the true time-varying parameter similar to (17) is also obtained when considering $(X_{t})_{t\in[n]}$-decorrelating partitions that are not orthogonal, i.e., such that (6) holds but the subspaces in the partition are not necessarily pairwise orthogonal. In this case, in particular, we can still find an invariant and residual parameter as in (14) and (16), respectively, where the matrix $U$ is now a non-orthogonal irreducible joint block diagonalizer.
@@
5:$U,\{S_{j}\}_{j=1}^{q}\leftarrow\operatorname{irreducibleJointBlockDiagonalizer}((\Sigma_{t})_{t\in[n]})$triangleright Prop. 1
@@
7:$S^{\operatorname{inv}},S^{\operatorname{res}}\leftarrow\emptyset$
8:for $j=1,\dots,q$ do
9:     $\Pi_{\mathcal{S}_{j}}\leftarrow U^{S_{j}}(U^{S_{j}})^{\top}$
10:     if $\Pi_{\mathcal{S}_{j}}\gamma_{0,t}$ constant in $[n]$  then $S^{\operatorname{inv}}\leftarrow S^{\operatorname{inv}}\cup S_{j}$triangleright see Prop. 2
@@
13:$\Pi_{\mathcal{S}^{\operatorname{inv}}}\leftarrow U^{\operatorname{inv}}(U^{\operatorname{inv}})^{\top}$
@@
[…]
Remark 2 (Changes of $\gamma_{0,t}$ and $\Sigma_{t}$ over time).
For the ISD procedure to work in practice, we assume that the covariance matrices $\Sigma_{t}$ are approximately constant in small time windows (e.g., smoothly changing or piece-wise constant). This is required to obtain accurate estimates of $\Sigma_{t}$, which we then (approximately) joint block diagonalize (see Section 4.1.1). Similarly, we assume that $\gamma_{0,t}$ is approximately constant in the adaptation window, in order to perform the adaptation step (see Setting 1). In the historical data, however, $\gamma_{0,t}$ can vary arbitrarily fast in the residual subspace ($\beta^{\operatorname{inv}}$ remains constant, of course). Intuitively, even if $\gamma_{0,t}$ does not change in a structured way, its projection $\Pi_{\mathcal{S}^{\operatorname{inv}}}\gamma_{0,t}$ on the invariant subspace remains constant across time points. We provide an example in which $\gamma_{0,t}$ is quickly varying in Appendix A.2.
3 Analysis of the two ISD tasks: zero-shot generalization and time adaptation
We now analyze the two prediction tasks that can be solved with the ISD framework introduced in the previous sections, namely (i) the zero-shot task and (ii) the adaptation task. We consider Setting 1 and assume for (i) that only the historical data is available, while for (ii) we assume to also have access to the adaptation data. Throughout the remainder of this section we assume that the invariant and residual subspaces are known (they can be computed from the joint distributions), which simplifies the theoretical analysis.